OpenStudy (anonymous):

Solve x^2 + 10x + 7 = 0 by completing the square. Which equation is used in the process? a. (x+ 10)^2= 93 b. (x+ 5)^2= 32 c. (x+ 5)^2= 18

OpenStudy (ash2326):

@bcheyennehaynes if we want to solve any quadratic equation by completing the square, we need to create the term (x+a)^2, here we have \[x^2+10x+7=0\] Can you notice the coefficient of x?

OpenStudy (anonymous):

(x+5)^2-18 = 0 (x+5)^2 = 18

OpenStudy (anonymous):

c